Two positive point charges a distance b apart have sum Q. For what values of the charges is the Coulomb force between them a maximum?
 
I have understood the question just not able to understand how to maximize q1q2, please tell me.

Grade:12

2 Answers

Vikas TU
4 years ago
Dear student 
For coulombic force to be maximum . 
Take dF/dt =0 
HHope this will clear your doubt .
Good Luck Cheers 
Arun
4 years ago
 

let parts are sepereted at a distance of d

so repulsion force F= k q(Q-q)/d2

     here Q and d are constant so for maximum force differentiate the force w.r.t q

 dF/dq = k/d[Q -2q]

 for maximum force dF/dq=0

                             Q-2q=0

Q/q = 2